OpenStudy (anonymous):

70% of what number is 224

OpenStudy (mitu12):

IS ---- OF 224 ---- 70 3.2 320 %...........ANSWER

OpenStudy (mitu12):

do you understand it

OpenStudy (triciaal):

@mitu12 no and second the answer has to be a number not % "...of what number" @Atiqullah let the number = x and remember % = per 100 then 70/100 * x = 224 solve for x x = 224 *100/70 = 320 70% of 320 is 224

jhonyy9 (jhonyy9):

try using the rule of 3 simples 70 % ----------- 224 100 % ---------- x ------------------------ x = (100*224)/70 = (10*224)/7 = 2240/7 = 320 so this number is 320 hope so much that this is understandably sure easy good luck